Project

General

Profile

Actions

Ops/Comm Scheduling meeting

Location : CR21 in R1 @ 15:30 pm on Tuesday 31st

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
  • DS readback current requires calibration (DL) : stands until close to powerup
  • Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
  • Document 'magnet off' buttons (Josef), when complete send to Andy & DC's. Must be a standalone document.
  • Compressor failure alarm for ISIS monitoring (Pierrick)
  • Update electrical work document (Steve G/Andy G)
  • DC report
  • Work plan for the next week

Ongoing work

  • SS QD/QP commissioning
    • DCCT readout on rampdown
    • Lakeshore studies (Pierrick, requires SC permit & possible shorting at link box)
  • Electrical work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
  • Water system work
    • Are we going to be ready for running by June?
  • FC cooldown
    • Magnet off button on FC? (Require consistency between magnets)
  • Target fibre work
  • ODH alarm issues
  • PRY movement calculations
  • Planning for July running
    • Proposed running plan
      • June 28 - July 3 : reserved for magnet work - we roster a single reserve shift
      • July 4 : expert startup
      • July 4 - July 9 : one 8-hour data-taking shift (2 people for 8 hours)
      • July 10 - July 16 : two 8-hour data-taking shifts ( 4 people for 16 hours)
      • July 17 - July 28 : Full 24 hour running
    • Run Preparation List
    • What needs to be ready to run?
      • SSH lockout procedure
      • Cooling water system
      • Controls and Monitoring including SSU/SSD/FC
      • Add items here...
    • BLOC training, 7th & 8th June
  • AOB
  • DONM : 7/6/2016 IN CR11 in R3

Minutes

Present : SB, HV, MP,

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
  • DS readback current requires calibration (DL) : stands until close to powerup / need people from Daresbury
  • Henry, Craig, Paul and Pierrick to look into the water detectors in the trench : Still being worked on. Stands
  • Document 'magnet off' buttons (Josef), when complete send to Andy & DC's. Must be a standalone document.
  • Compressor failure alarm for ISIS monitoring (Pierrick) : (stands)
  • Update electrical work document (Steve G/Andy G) : (stands)
  • DC report
    • See below
  • Work plan for the next week
    • Not up yet because require feedback from this meeting
    • Next Monday fire office is coming to look at emergency procedure for the H2 system (Monday @ 12:30 pm)

Ongoing work

  • SS QD/QP commissioning
    • DCCT readout on rampdown
    • Lakeshore studies (Pierrick, requires SC permit & possible shorting at link box)
      • SSU pretty much finished. On SSD two "missing" voltage taps have been found to be not-missing. SSD tuned up to the same level of QP as SSU.
      • DCCTs should show exactly what current is going through the magnets, however they are on the wrong path now. Pierrick has had a meeting
        with Trevor and he is going to look into this next week (Tues/Wed next week)
      • Lakeshores are better but still far from perfect (Pierrick) - needs some time reserved to work on them. Nominally Tuesday 7th morning / Wednesday afternoon reserved for Pierrick's work on Lakeshore.
      • John G : what about the heaters. Pierrick : need to get in contact with Gallagher. Heaters are connected but not switched on. We need better
        blowers. Action : get in touch with Andy Gallager ASAP.
  • Electrical work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx : Action Item : Need new plan (Andy G / Steve G)
  • Water system work
    • Chiller is connected and chiller is full of water. Permit has not been signed off by electrical officer yet (hopefully tomorrow). Once we
      get that we'll run it and will move over to compressors if that's what we want to do. Are we going to start work on the cooling system?
      Hoping to have the chiller up and running by end of the week. When do we break in to the cooling system? Colin - not until after the ISIS
      run. John agrees.
    • Colin : We need a plan for what the US team wants to do well in advance of when they arrive. *Action Item : * DC to press the US team for a work plan.
    • Are we going to be ready for running by June? Water won't go in until after the ISIS run. Josef : what happens if the water fails in the
      meantime? We must have a document in case this happens. Action Item : ( Josef, John, DC, Pierrick ) to write a document on what to do if the
      compressors or the water fail. Henry : can we test the conventional magnets next week. We'll test this on Thursday nominally.
  • FC cooldown
    • 150K on the coldmass at the rate of 12-13 K per day. Another 10 days and it's cold.
    • Pierrick : what still needs to be done on the controls etc ; Josef : Daresbury still needs to do some connections and check the system. Once they have signed off Josef, John C and Pierrick will run it and check it. Then we'll start ramping it up in stages. This may take a week or two.
      Ideally would like to ramp in conjunction with SSU. Pierrick : are we going to operate in solenoid or flip mode? Chris : plan to run in solenoid
      mode @ 3 T. Action : Chris to tell us if we want to go about 3 T. Chris: We have a pragmatic set of currents for the next user run for
      beam commissioning. If we're happy with the way things are going after that we can think about moving to 4 T, but that is a decision for later
      user cycles.
  • Decision : * Plan to run this user cycle at 3T. Can revisit this if all goes well.
    • Magnet off button on FC? (Require consistency between magnets)
  • Target fibre work
    • Proceeding as expected.
  • ODH alarm issues
    • ODH alarms going off - seems to be due to the age of the sensors. They will be replaced on 13/14 June.
  • PRY movement calculations
    • When SSU was powered on the weekend. There were indications that the PRY was moving as the magnet was ramping. After SSU was ramped
      down the East Wall had moved 0.22 mm. Colin : Steve Plate's calculations had 1 mm of movement at full current. Jason will work on further
      calculations based on new magnet configurations next week. Pierrick : request that a monitor is put on the top plate as well as the bottom
      plate. Action Item : get a record of the measurements of movements that have been made to the engineers to let them think about it. Victoria : can
      we power it and make dedicated measurements. Colin: yes we can do that. Action Item : Victoria/Pierrick to discuss dedicated movement measurements.
      Josef : should we ask Holge to looking to calculating the forces? Holge : in general 1 mm movement of PRY is to be expected. Steve's model
      doesn't include any movement which comes from tiedowns to the floor so Steve's model is incomplete. In general mm scale measurement is to
      be expected. Action Item : check that this information is in the archiver (Victoria). Action Item : we should also torque the bolts and verify the sensors are operating as expected.
  • Outreach : June 9th
  • Planning for July running
    • Proposed running plan
      • June 28 - July 3 : reserved for magnet work - we roster a single reserve shift
      • July 4 : expert startup : Action : set up a doodle poll because I forgot about Neutrino 2016 damn it.
      • July 4 - July 9 : one 8-hour data-taking shift (2 people for 8 hours)
      • July 10 - July 16 : two 8-hour data-taking shifts ( 4 people for 16 hours)
      • July 17 - July 28 : Full 24 hour running
    • Run Preparation List
    • What needs to be ready to run?
      • SSH lockout procedure
      • Cooling water system
      • Controls and Monitoring including SSU/SSD/FC
      • Add items here...
    • BLOC training, 7th & 8th June
  • AOB : Smug cat

Updated by Boyd, Steven about 6 years ago ยท 10 revisions